OpenStudy (learner31):

costheta = (1- tan^2(theta/2))/ (1 + tan^2(theta/2))

OpenStudy (daniel.ohearn1):

You want to find theta?

OpenStudy (daniel.ohearn1):

One thing you can do think about the pytagorean identities first off

OpenStudy (daniel.ohearn1):

s^2 + c^2 = 1 divide each side by cosine we get tan^2 + 1 = sec^2